Problem 30.1 A 2.08 m -ong coil containing 225 loops is wound on an iron core (average u 1850uo) along with a second coil of 115 loops. The loops of each coil have a radius of 1.10 cm If the current in the first coil drops uniformly from 12.0 A to zero in 96.0 ms. Part A Determine the mutual inductance M. Submit My Answers Give Up Part B Determine the emf induced in the second coi e2 Submit My Answers Give Up EMI Provide F Continue eedback
Answer #1

Area of the coil A = π×(0.011)2 4-3.8×10-4 m2 Value of μ = 1850×4m x 10-7 μ = 2.32 x 10-3 Part A Mutual Inductance of the coiPart B emf induced in the second coil dI dt є -001098-12 E0.0109 96x10% 1.3625 V